# Improved Approximation Algorithms for Chromatic and Pseudometric-Weighted Correlation Clustering

### v1

@@ Appendix B Proof for Theorem 2
I $+$ 00 $x$ 00 $y$
-- $1$ $1-x$ $1$ $1-y$
II $+$ $\frac{5}{3}z$ $x$ $\frac{5}{3}z$ $y$
-- $1-\frac{5}{3}z$ $1-x$ $1-\frac{5}{3}z$ $1-y$
III $+$ $1$ $x$ $1$ $y$
-- 00 $1-x$ 00 $1-y$
IV $+$ $\frac{5}{3}y+\frac{5}{3}z-\frac{50}{9}yz$ $(1-\frac{25}{9}yz)x$ $\frac{5}{3}z$ $y$
-- $(1-\frac{5}{3}y)(1-\frac{5}{3}z)$ $(1-\frac{25}{9}yz)(1-x)$ $1-\frac{5}{3}z$ $1-y$
V $+$ $1-\frac{5}{3}y$ $(1-\frac{5}{3}y)x$ $1$ $y$
-- 00 $(1-\frac{5}{3}y)(1-x)$ 00 $1-y$
VI $+$ 00 00 $1$ $y$
-- 00 00 00 $1-y$
VII $+$ $1-\frac{5}{3}y$ $(1-\frac{5}{3}y)x$ $1-\frac{5}{3}x$ $(1-\frac{5}{3}x)y$
-- 00 $(1-\frac{5}{3}y)(1-x)$ 00 $(1-\frac{5}{3}x)(1-y)$
@@ VII $+$ $1-\frac{5}{3}y$ $(1-\frac{5}{3}y)x$ $1-\frac{5}{3}x$ $(1-\frac{5}{3}x)y$
I $+$ 00 $z$
-- $1$ $1-z$
II $+$ 00 $z$
-- $1$ $1-z$
III $+$ 00 $z$
-- $1$ $1-z$
IV $+$ $\frac{5}{3}y$ $z$
-- $1-\frac{5}{3}y$ $1-z$
V $+$ $\frac{5}{3}y$ $z$
--$1-\frac{5}{3}y$ $1-z$
VI $+$ $1$ $z$
-- 00 $1-z$
VII $+$ $\frac{5}{3}x+\frac{5}{3}y-\frac{50}{9}xy$ $(1-\frac{25}{9}xy)z$
@@
B.1 Triangle inequality is tight.
B.1.1 Region I, II, III, VI
Affine formulae; refer to Table 5. Also note that the factor $10/3$ is tight for $(0,\frac{2}{5},\frac{2}{5})$ with (+,+,+)(+,+,+)( + , + , + ),
@@ I $(0,0)$ $+$ 00 00 00
$(0,\frac{2}{5})$ $+$ 00 $4/3$ $4/3$
-- $7/3$ $1$ $1$
$(\frac{1}{5},\frac{1}{5})$ $+$ $2/3$ $2/3$ $4/3$
-- $5/3$ $5/3$ $1$
II $(0,\frac{2}{5})$ $+$ $-2/3$ $2/3$ $4/3$
-- $3$ $5/3$ $1$
$(\frac{1}{5},\frac{1}{5})$ $+$ 00 00 $4/3$
-- $7/3$ $7/3$ $1$
$(\frac{1}{5},\frac{2}{5})$ $+$ $-1/3$ $1/3$ $2$
-- $8/3$ $2$ $1/3$
$(\frac{3}{10},\frac{3}{10})$ $+$ 00 00 $2$
@@ II $(0,\frac{2}{5})$ $+$ $-2/3$ $2/3$ $4/3$
III $(\frac{1}{5},\frac{2}{5})$ $+$ $-1/3$ $1/3$ $2$
-- $8/3$ $2$ $1/3$
$(\frac{3}{10},\frac{3}{10})$ $+$ 00 00 $2$
@@ III $(\frac{1}{5},\frac{2}{5})$ $+$ $-1/3$ $1/3$ $2$
$(\frac{2}{5},\frac{2}{5})$ $+$ $1/3$ $1/3$ $8/3$
-- $2$ $2$ $-1/3$
VI $(0,\frac{3}{5})$ $+$ 00 $1$ $1$
-- 00 $4/3$ $4/3$
$(0,1)$ $+$ 00 $7/3$ $7/3$
-- 00 00 00
$(\frac{2}{5},\frac{3}{5})$ $+$ 00 $1$ $7/3$
-- 00 $4/3$ 00
B.1.2 Region IV
The argument here is analogous with A.1.2: $(1-\frac{25}{9}yz)x\leq(1-\frac{25}{9}yz)(1-x)$ as $x\leq\frac{1}{2}$; $\frac{5}{3}y+\frac{5}{3}z-\frac{50}{9}yz\geq(1-\frac{5}{3}y)(1-\frac{5}{3}z)$ as
@@ B.1.2 Region IV
and $y,z\in\left[2/5,3/5\right]$. Therefore, $\frac{10}{3}\cdot e.lp_{w}-e.cost_{w}$ is smaller with sign ‘$+$’.
Replacing $x=z-y$,
@@
Fixing $y$, this formula is concave of $z$, hence minimized in either $z=y$ or $z=3/5$, which yield $\frac{50}{9}y^{2}-\frac{10}{3}y$ and $\frac{50}{9}y^{2}-5y+1$ each. Within the range $y\in[2/5,3/5]$, the minimum value of each is $-1/9$ in $y=2/5$ and $-1/8$ in $y=9/20$. Thus, $\frac{10}{3}\cdot e.lp_{w}-e.cost_{w}\geq-1/8$.
Finally, $\frac{10}{3}\cdot e.lp_{u}-e.cost_{y}\geq 1/3$, minimized at $(y,z)=(2/5,3/5)$ with ‘$+$’ sign; $\frac{10}{3}\cdot e.lp_{v}-e.cost_{v}\geq 2/3$, minimized at $(y,z)=(2/5,2/5)$ with ‘$+$’ sign.
@@
B.1.3 Region V
The argument here is analogous with A.1.3: $\frac{10}{3}\cdot e.lp_{w}-e.cost_{w}$ is smaller with sign ‘$+$’, which is $(\frac{10}{3}x-1)(1-\frac{5}{3}y)$. This has a minimum value of $-1/9$ in $(x,y)=(1/5,2/5)$.
The formula for $\frac{10}{3}\cdot e.lp_{v}-e.cost_{v}$ is either $5/3(2x+y)$ or $7/3-5/3(2x+y)$, whose global minimum value is 00 in $(x,y)=(2/5,3/5)$ with sign ‘--’.
The formula for $\frac{10}{3}\cdot e.lp_{u}-e.cost_{u}$ is either $110/3\cdot y-$ 1 or $10/3(1-y)$, whose global minimum value is $1/3$ in $y=2/5$ with sign ‘$+$’. Thus, $(\frac{10}{3}\cdot e.lp_{w}-e.cost_{w})+(\frac{10}{3}\cdot e.lp_{u}-e.cost_{u})$ and $(\frac{10}{3}\cdot e.lp_{u}-e.cost_{u})+(\frac{10}{3}\cdot e.lp_{v}-e.cost_{v})$ are at least $-1/9+1/3=2/9\geq 0$.
@@ B.1.3 Region V
B.1.4 Region VII
The argument here is analogous with A.1.4: $\frac{10}{3}\cdot e.lp_{w}-e.cost_{w}$ is smaller with sign ‘$+$’, which is $(\frac{10}{3}x-1)(1-\frac{5}{3}y)$. This has a minimum value of 00 in $(x,y)=(0.4,0.6)$.
The formula for $\frac{10}{3}\cdot e.lp_{u}-e.cost_{u}$ is either $(1-\frac{5}{3}x)(\frac{10}{3}y-1)$ or $\frac{10}{3}(1-\frac{5}{3}x)(1-y)$, whose global minimum value is $1/9$ in $(x,y)=(2/5,2/5),(1/2,1/2)$ with sign ‘$+$’.
@@ B.1.4 Region VII
since $z\geq\frac{4}{5}$. Therefore, the formula with sign ‘--’ is smaller, which is $\frac{10}{3}(1-\frac{25}{9}xy)(1-z)-(1-\frac{5}{3}x)(1-\frac{5}{3}y)$.
@@
This function decreases in the range $z\in[4/5,1]$, hence the minimum value is $-1/36$ in $z=1$. Therefore, $(\frac{10}{3}\cdot e.lp_{w}-e.cost_{w})+(\frac{10}{3}\cdot e.lp_{u}-e.cost_{u})$ and $(\frac{10}{3}\cdot e.lp_{u}-e.cost_{u})+(\frac{10}{3}\cdot e.lp_{v}-e.cost_{v})$ are at least $1/9-1/36=1/12\geq 0$.
@@
B.2 All coordinates belong to endpoint of segments.
@@
$(\frac{3}{5},\frac{3}{5},1)$ $+$ 00 00 00
-- 00 00 00
$(\frac{2}{5}-\delta,1,1)$ $+$ 00 $7/3$ $7/3$
-- 00 00 00
$(\frac{2}{5}+\delta,1,1)$ $+$ 00 $7/9$ $7/9$
-- 00 00 00
$(\frac{3}{5},1,1)$ $+$ 00 00 00
-- 00 00 00
$(1,1,1)$ $+$ 00 00 00
-- 00 00 00

### v2

@@ Appendix A Proof for Theorem 2
I ++ 0 $x$ 0 $y$
- 11 1x1-x 11 1y1-y
II ++ $\frac{5}{3}z$ $x$ $\frac{5}{3}z$ $y$
- $1-\frac{5}{3}z$ 1x1-x $1-\frac{5}{3}z$ 1y1-y
III ++ 11 $x$ 11 $y$
- 0 1x1-x 0 1y1-y
IV ++ $\frac{5}{3}y+\frac{5}{3}z-\frac{50}{9}yz$ $(1-\frac{25}{9}yz)x$ $\frac{5}{3}z$ $y$
- $(1-\frac{5}{3}y)(1-\frac{5}{3}z)$ $(1-\frac{25}{9}yz)(1-x)$ $1-\frac{5}{3}z$ 1y1-y
V ++ $1-\frac{5}{3}y$ $(1-\frac{5}{3}y)x$ 11 $y$
- 0 $(1-\frac{5}{3}y)(1-x)$ 0 1y1-y
VI ++ 0 0 11 $y$
- 0 0 0 1y1-y
VII ++ $1-\frac{5}{3}y$ $(1-\frac{5}{3}y)x$ $1-\frac{5}{3}x$ $(1-\frac{5}{3}x)y$
- 0 $(1-\frac{5}{3}y)(1-x)$ 0 $(1-\frac{5}{3}x)(1-y)$
@@
I ++ 0 $z$
- 11 1z1-z
II ++ 0 $z$
- 11 1z1-z
III ++ 0 $z$
- 11 1z1-z
IV ++ $\frac{5}{3}y$ $z$
- $1-\frac{5}{3}y$ 1z1-z
V ++ $\frac{5}{3}y$ $z$
-$1-\frac{5}{3}y$ 1z1-z
VI ++ 11 $z$
- 0 1z1-z
VII ++ $\frac{5}{3}x+\frac{5}{3}y-\frac{50}{9}xy$ $(1-\frac{25}{9}xy)z$
@@
A.1 Triangle inequality is tight.
A.1.1 Region I, II, III, VI
Affine formulae; refer to Table 5. Also note that the factor $10/3$ is tight for $(0,\frac{2}{5},\frac{2}{5})$ with $(+,+,+)$,
@@ I $(0,0)$ ++ 0 0 0
$(0,\frac{2}{5})$ ++ 0 $4/3$ $4/3$
- $7/3$ 11 11
$(\frac{1}{5},\frac{1}{5})$ ++ $2/3$ $2/3$ $4/3$
- $5/3$ $5/3$ 11
II $(0,\frac{2}{5})$ ++ 2/3-2/3 $2/3$ $4/3$
- 33 $5/3$ 11
$(\frac{1}{5},\frac{1}{5})$ ++ 0 0 $4/3$
- $7/3$ $7/3$ 11
$(\frac{1}{5},\frac{2}{5})$ ++ 1/3-1/3 $1/3$ 22
- $8/3$ 22 $1/3$
$(\frac{3}{10},\frac{3}{10})$ ++ 0 0 22
@@ II $(0,\frac{2}{5})$ ++ 2/3-2/3 $2/3$ $4/3$
III $(\frac{1}{5},\frac{2}{5})$ ++ 1/3-1/3 $1/3$ 22
- $8/3$ 22 $1/3$
$(\frac{3}{10},\frac{3}{10})$ ++ 0 0 22
@@ III $(\frac{1}{5},\frac{2}{5})$ ++ 1/3-1/3 $1/3$ 22
$(\frac{2}{5},\frac{2}{5})$ ++ $1/3$ $1/3$ $8/3$
- 22 22 1/3-1/3
VI $(0,\frac{3}{5})$ ++ 0 11 11
- 0 $4/3$ $4/3$
$(0,1)$ ++ 0 $7/3$ $7/3$
- 0 0 0
$(\frac{2}{5},\frac{3}{5})$ ++ 0 11 $7/3$
- 0 $4/3$ 0
A.1.2 Region IV
The argument here is analogous with 4.4: $(1-\frac{25}{9}yz)x\leq(1-\frac{25}{9}yz)(1-x)$ as $x\leq\frac{1}{2}$; $\frac{5}{3}y+\frac{5}{3}z-\frac{50}{9}yz\geq(1-\frac{5}{3}y)(1-\frac{5}{3}z)$ as
@@ A.1.2 Region IV
and $y,z\in\left[2/5,3/5\right]$. Therefore, $\frac{10}{3}\cdot e.lp_{w}-e.cost_{w}$ is smaller with sign ‘++’.
Replacing x=zyx=z-y,
@@
Fixing $y$, this formula is concave of $z$, hence minimized in either $z=y$ or $z=3/5$, which yield $\frac{50}{9}y^{2}-\frac{10}{3}y$ and $\frac{50}{9}y^{2}-5y+1$ each. Within the range $y\in[2/5,3/5]$, the minimum value of each is 1/9-1/9 in $y=2/5$ and 1/8-1/8 in $y=9/20$. Thus, $\frac{10}{3}\cdot e.lp_{w}-e.cost_{w}\geq-1/8$.
Finally, $\frac{10}{3}\cdot e.lp_{u}-e.cost_{y}\geq 1/3$, minimized at $(y,z)=(2/5,3/5)$ with ‘++’ sign; $\frac{10}{3}\cdot e.lp_{v}-e.cost_{v}\geq 2/3$, minimized at $(y,z)=(2/5,2/5)$ with ‘++’ sign.
@@
A.1.3 Region V
The argument here is analogous with 4.4: $\frac{10}{3}\cdot e.lp_{w}-e.cost_{w}$ is smaller with sign ‘++’, which is $(\frac{10}{3}x-1)(1-\frac{5}{3}y)$. This has a minimum value of 1/9-1/9 in $(x,y)=(1/5,2/5)$.
The formula for $\frac{10}{3}\cdot e.lp_{v}-e.cost_{v}$ is either $5/3(2x+y)$ or 7/35/3(2x+y)7/3-5/3(2x+y), whose global minimum value is 0 in $(x,y)=(2/5,3/5)$ with sign ‘-’.
The formula for $\frac{10}{3}\cdot e.lp_{u}-e.cost_{u}$ is either $10/3\cdot y-1$ or 10/3(1y)10/3(1-y), whose global minimum value is $1/3$ in $y=2/5$ with sign ‘++’. Thus, $(\frac{10}{3}\cdot e.lp_{w}-e.cost_{w})+(\frac{10}{3}\cdot e.lp_{u}-e.cost_{u})$ and $(\frac{10}{3}\cdot e.lp_{u}-e.cost_{u})+(\frac{10}{3}\cdot e.lp_{v}-e.cost_{v})$ are at least $1/9+1/3=2/9\geq 0$.
@@ A.1.3 Region V
A.1.4 Region VII
The argument here is analogous with 4.4: $\frac{10}{3}\cdot e.lp_{w}-e.cost_{w}$ is smaller with sign ‘++’, which is $(\frac{10}{3}x-1)(1-\frac{5}{3}y)$. This has a minimum value of 0 in $(x,y)=(0.4,0.6)$.
The formula for $\frac{10}{3}\cdot e.lp_{u}-e.cost_{u}$ is either $(1-\frac{5}{3}x)(\frac{10}{3}y-1)$ or $\frac{10}{3}(1-\frac{5}{3}x)(1-y)$, whose global minimum value is $1/9$ in $(x,y)=(2/5,2/5),(1/2,1/2)$ with sign ‘++’.
@@ A.1.4 Region VII
since $z\geq\frac{4}{5}$. Therefore, the formula with sign ‘-’ is smaller, which is $\frac{10}{3}(1-\frac{25}{9}xy)(1-z)-(1-\frac{5}{3}x)(1-\frac{5}{3}y)$.
@@
This function decreases in the range $z\in[4/5,1]$, hence the minimum value is 1/36-1/36 in $z=1$. Therefore, $(\frac{10}{3}\cdot e.lp_{w}-e.cost_{w})+(\frac{10}{3}\cdot e.lp_{u}-e.cost_{u})$ and $(\frac{10}{3}\cdot e.lp_{u}-e.cost_{u})+(\frac{10}{3}\cdot e.lp_{v}-e.cost_{v})$ are at least $1/9-1/36=1/12\geq 0$.
@@
A.2 All coordinates belong to endpoint of segments.
@@
$(\frac{3}{5},\frac{3}{5},1)$ ++ 0 0 0
- 0 0 0
$(\frac{2}{5}-\delta,1,1)$ ++ 0 $7/3$ $7/3$
- 0 0 0
$(\frac{2}{5}+\delta,1,1)$ ++ 0 $7/9$ $7/9$
- 0 0 0
$(\frac{3}{5},1,1)$ ++ 0 0 0
- 0 0 0
$(1,1,1)$ ++ 0 0 0
- 0 0 0